A monopoly firm faces a linear demand curve P = 10 – 0.5 Q. Its MC is constant at Rs. 4. 

What is the welfare loss on account of monopoly?


1
Expert's answer
2022-02-22T13:03:54-0500

"P = 10 \u2013 0.5 Q."

"TR= P.Q= Q(10-0.5Q)"

"= 10Q-0.5Q^2"

"MR= 10- Q"

Under Monopoly,

"MR= MC"

"10-Q= 4"

"Q= 6"

"P= 10- 0.5(6)= 7"

Under perfect competitive market

MR= P

"10-0.5Q= 4"

"0.5Q= 6"

Q= 12

"P= 10- 0.5(12)= 4"

Welfare loss"= \frac{1}{2}\times (P_1-P_2)\times(Q_1-Q_2)"

"= \frac{1}{2}\times (7-4)\times(12-6)"

"= \frac{1}{2}\times 3\times 6= 6"